OpenStudy (anonymous):

For 0°≤θ<360°, how many roots does the equation tanθ+2sinθ=0 have?

jimthompson5910 (jim_thompson5910):

tan(x) + 2*sin(x) = 0 sin(x)/cos(x) + 2*sin(x) = 0 sin(x)( 1/cos(x) + 2 ) = 0 sin(x) = 0 or 1/cos(x) + 2 = 0 I'll let you finish up

OpenStudy (anonymous):

is it 3?

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

sinx=0 has 2 roots right?

OpenStudy (anonymous):

cosx=-1/2..... 4?

jimthompson5910 (jim_thompson5910):

yes if sin(x) = 0, then x = 0 or x = 180

jimthompson5910 (jim_thompson5910):

1/cos(x) + 2 = 0 1/cos(x) = -2 cos(x) = -1/2 x = 120 or x = 240

jimthompson5910 (jim_thompson5910):

So there are 4 solutions in the interval 0 <= x < 360

OpenStudy (anonymous):

ohh, yes, thank you

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

and which of the following equations has/have solutions? A. 2cos^(2)θ-sin^(2)θ=1 B. 2cos^(2)θ-sin^(2)θ=2 C. 2cos^(2)θ-sin^(2)θ=3

jimthompson5910 (jim_thompson5910):

Your best bet is to graph each of them if the equation has x-intercepts, then the equation has solutions